2. Income taxes

The following table shows income tax rates in Econoland.

Annual Income

Tax Rate

Up to $60,000 10%
From $60,001 to $90,000 20%
Over $90,000 30%

This is an example of a_____ income tax.

Van, a resident of Econoland, currently works 20 hours a week and earns an annual income of $60,000. After paying income taxes, Van receives ______ per year.

If Van works an additional 10 hours a week (30 hours a week total), his annual income will be $90,000. After paying income taxes, Van will receive______ per year. In other words, by working an additional 10 hours per week, Van will receive an additional______ .

If Van works 10 more hours in a week (40 hours a week total), his annual income will further increase to $120,000. After paying income taxes, Van will receive______ per year. In other words, by working 10 more hours per week, Van will receive_______ more than what he would receive if he works only 30 hours a week.

As Van works more, the additional amount he receives from working 10 more hours_____ , which means his cost of leisure______.   .
Answer #1

Ans: This is an example of a progressive  income tax.

Ans: Van, a resident of Econoland, currently works 20 hours a week and earns an annual income of $60,000. After paying income taxes, Van receives $54000 per year.

Explanation:

Total tax amount = $60,000 * 0.10 = $6000

Disposable income = $60000 - $6000 = $54000

Ans: If Van works an additional 10 hours a week (30 hours a week total), his annual income will be $90,000. After paying income taxes, Van will receive $78000 per year. In other words, by working an additional 10 hours per week, Van will receive an additional $24000.

Explanation:

Total tax amount up to $60000= $60,000 * 0.10 = $6000

Total tax amount from $60001 to $90000 = $30000 * 0.20 = $6000

Disposable income = $90000 - ( $6000 + $6000) = $78000

By working an additional 10 hours per week, Van will receive an additional = $78000 - $54000 = $24000

Ans: If Van works 10 more hours in a week (40 hours a week total), his annual income will further increase to $120,000. After paying income taxes, Van will receive $99000 per year. In other words, by working 10 more hours per week, Van will receive $21000 more than what he would receive if he works only 30 hours a week.

Explanation:

Total tax amount up to $60000= $60,000 * 0.10 = $6000

Total tax amount from $60001 to $90000 = $30000 * 0.20 = $6000

Total tax amount from $90001 to $120000 = $30000 * 0.30 = $9000

Disposable income = $120000 - ( $6000 + $6000 +$9000) = $99000

By working an additional 10 hours per week, Van will receive an additional = $99000 - $78000 = $21000

Ans: As Van works more, the additional amount he receives from working 10 more hours increases , which means his cost of leisure increases.